Two points A and B have coordinates (3,6) and (-3,0) respectively.

Obtain the equation of the circle for which the line segment AB is the diameter


1
Expert's answer
2021-09-09T00:40:39-0400

The general equation for a circle is "( x - h )^2 + ( y - k )^2 = r^2," where ("h, k )" is the center and "r" is the radius.

The line segment "AB" is the diameter. Then the center of the circle "C(h, k)" is the middle of the line segment "AB."


"h=\dfrac{x_A+x_B}{2}=\dfrac{3-3}{2}=0"

"k=\dfrac{y_A+y_B}{2}=\dfrac{6+0}{2}=3"

Center "C(0,3)."


The radius of the circle is half of the diameter


"r=\dfrac{\sqrt{(x_B-x_A)^2+(y_B-y_A)^2}}{2}"

"=\dfrac{\sqrt{(-3-3)^2+(0-6)^2}}{2}=3\sqrt{2}"

Substitute


"( x - 0 )^2 + ( y - 3 )^2 = (3\sqrt{2})^2"



The general equation of the circle  for which the line segment "AB" is the diameter is 


"x^2 + ( y - 3 )^2 = 18"